OpenStudy (anonymous):

f(x)=x^3+3e^x (a) What is the derivative of f at x=0? (b) Using the (correct) value of f′(0), give the equation of the tangent line to the curve y=f(x) at the point (0,f(0))=(0,3). Your answer must be in the form of an equation for y in terms of x, that is, of the form y=L(x) for some linear function L(x).

OpenStudy (anonymous):

@abb0t

OpenStudy (amistre64):

fairly simple derivative, what do you come up with?

OpenStudy (anonymous):

3x^2 + 3e^x

OpenStudy (amistre64):

good, and when x=0?

OpenStudy (anonymous):

undefine?

OpenStudy (anonymous):

cause x can never be 0

OpenStudy (amistre64):

what property are you trying to distort to get that?

OpenStudy (anonymous):

3(x^2 +e^x) since e^x can't be 0 then it would be undefine right?

OpenStudy (amistre64):

let x=0 3(0^2 + e^0) we are replacing x with 0, not e^x with 0

OpenStudy (anonymous):

so the answer for a is 3?

OpenStudy (amistre64):

yes, the derivative (the slope at x=0) is 3

OpenStudy (amistre64):

now, when we know a slope and a point, how do we construct a line?/

OpenStudy (anonymous):

y=mx+b

OpenStudy (amistre64):

thats slope and intercept .... which would work here if you recognize the given point (0,3) as a y intercept

OpenStudy (amistre64):

if you can tell if its an intercept, we could just as well use the point slope format \[y-y_o=m(x-x_0)\]

OpenStudy (amistre64):

*if you cant tell ...

OpenStudy (anonymous):

i dunno then

OpenStudy (amistre64):

if you dunno if its an intercept, lets use the point slope form to work with

OpenStudy (anonymous):

it's not an intercept that's for sure

OpenStudy (amistre64):

for a slope of 3, and a point (0,3) y - 3 = 3(x - 0) now we can work it into a y=_______ construction

OpenStudy (anonymous):

y=3x+3

OpenStudy (amistre64):

it may not be sure for you, but its pretty sure to me that it is a y intercept ..... but thats not important at the moment. we simply need to use it in some useful fashion and a point slope format works either way

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

i just don't get the slope and the point part

OpenStudy (amistre64):

what does the derivative of a function tell us about the function?

OpenStudy (anonymous):

the slope

OpenStudy (amistre64):

and, they stated the point they wanted right? (0,3)

OpenStudy (amistre64):

so all it boils down to is using those parts to make the line equation for the tangent to the curve

OpenStudy (anonymous):

oh and then you just solve for yfinal okay i get it

OpenStudy (amistre64):

practice makes it almost second nature

OpenStudy (anonymous):

okay thanks!